if n+L =6, then total number of subshells would be

if  n+L =6, then total number of subshells would be

Grade:Select Grade

1 Answers

Aarti Gupta
8 years ago
The value of n + l = 6 then,
n + l = 6
6 + 0 = 6
5 + 1 = 6 and
4 + 2 = 6
All above can be possible as values of ‘n’ are greater than the values of ‘l’
But,
3 + 3 = 6
2 + 4 = 6
1 + 5 = 6
0 + 6 = 6
3+ 3 not possible as value of l = (n--1),and other values cannot be possible as value of ‘n’ are less than l.
Hence total number of sub-shells would be 3.